# Multi-objective Combinatorial Methodology for Nuclear Reactor Site Assessment: A Case Study for the United States

### v1

@@ 2.1 Socioeconomic Characteristics
We extracted nuclear restrictions in each statesidering it is an uncontrollable parameter for a stakeholder, we define a moratorium as a deal breaker. We express all other restrictions in a single variable as the number of nuclear restrictions in the state. These five remaining restrictions are state legislature approval, state commissioner of environmental protection approval, proving that the construction of a nuclear facility will be economically feasible for ratepayers, demonstrable technology or a means for high-level waste disposal or reprocessing, and proving that the radioactive waste material disposal method will be safe.
We define state electricity price as the annual average electricity price in cents/kWh and electricity imports as the annual amount of electrical power imported by the state in millions of kWh. Our ”population sentiment towards nuclear energy” parameter is an average of the 10-year poll findings of sentiment values of the counties in any 20-mile radius We set state nuclear inclusive policy as a binary value that shows if the state administration has any decisions to support the development of nuclear energy, such as Renewable Portfolio Standards (RPSs), Renewable Portfolio Goals (RPGs), and Clean Energy Standards (CESs).
@@ 2.2 Safety Characteristics
We selected safety objectives characteristics that support siting a 1000MWt or higher power reactor [19]. Most of the safety objectives (1, 3, 4, 6, 7) are intersections with dangerous areas explained in the list of objectives provided earlier in this section. NRC Regulatory Guide 4.7, “General Site Suitability Criteria for Nuclear Power Stations” describes that siting nuclear reactors close to the publicly used lands depends on local jurisdictions and can be refused. For this reason, the protected land objective checks if the site intersects with American Indian reservations, correctional facilities, critical habitats, forests, hospitals, national monuments, national or state parks, schools or colleges, wild and scenic rivers, wilderness areas, or wildlife refuges. If the site intersects with any of these, the number of intersections is given.
@@ 2.3 Proximity Characteristics
Proximity objectives have been used as binary values as suggested by the STAND tool and previous research conducted in the area. For example, the 10 CFR 100 Regulatory Guide of the U.S. Nuclear Regulatory Commission (NRC) shows that at least a four-mile distance is required between the NPP and the population centers with 25,000 residents [27, 28]. In the STAND tool, this objective is set to 0 if the closest population center distance is closer than four miles. For centers with more than four miles, the objective is linearly scaled between [0,1], where the center with the highest distance from the plant is set to 1.
@@
The retiring power station objective shows if any coal, natural gas, or nuclear power station is in a one-mile radius in the next 20 years. For the sites with a retiring power station nearby, the distances are linearly scaled to the [0,1] range where the lowest-distance site is set to 1. The substation distance objective is the site distance to the closest electricity grid substation in miles, scaled to [0,1] range where the closest site to the substation is set to 1. We define transportation systems as major roads in the U.S.. The objective is set to 1 if there is a major road closer than 1 mile. For the sites that do not have a major road in a 1-mile radius, the values are linearly scaled to the [0,1] range where the lowest-distance site is set to 1. The streamflow objective shows if there are any freshwater supplies in 20-mile radius, with the minimum 50,000 gallons per minute (GPM) streamflow as recommended in the STAND tool.
The distance values in this section are the distance values that the STAND tool uses for thresholding and linearly scaling the attributes of the selected sites. We selected the generator retirement time (20 years), transportation method (major roads), and streamflow GPM value (50K) with the guidance of the STAND tool developers.
Lastly, an example of the created preprocessed dataset is given in Figure 1. Since the dataset is scaled, higher values for each attribute indicate favorable objectives. For example, 0.9 in the electricity price column indicates a low electricity price in the state, while 0.1 indicates a high electricity price.
@@ 2.4 Coal Power Plant Sites
According to the U.S. Department of Energy, there are currently 157 inactive and 237 active candidate coal power plant locations that are suitable for housing NPPs [10]. These sites are accompanied by pertinent data, including the 22 objectives outlined in the STAND framework, which are detailed in previous sections and in Table 1. However, the CPP dataset lacks the state federal information processing standard (FIPS) number and county FIPS number from its input data [^1] https://transition.fcc.gov/oet/info/maps/census/fips/fips.txt. Previous research identified 265 U.S. coal power plant sites as relevant for NPP siting [19]. Additionally, the coal-to-nuclear transition dataset was updated in October 2024 by the Fastest Path to Zero Initiative (FPTZ) Initiative at the University of Michigan. We have adapted the coal power plant dataset, which was previously examined using the weighted sum method, for multi-objective optimization. By using this dataset, we facilitated a one-to-one comparative analysis with the brownfield sites.
The coal-to-nuclear siting dataset is considerably smaller than the brownfield siting dataset, resulting in significantly shorter processing times. Unfortunately, the limited size also restricts the ability to train machine learning models. Consequently, we exclusively used the brownfield dataset to train the machine learning models. The CPP locations were processed separately, allowing for a comparison between the optimal locations identified in both datasets.
@@ 3 Research Methodology
This section describes the methodology used for the assessment of the brownfield and Coal sites described in Section 2, which consists of two major parts. First, we describe our high-dimensional non-dominated combinatorial search approach that is used to rank all sites based on how many times they show up in the Pareto-front when considering every possible combination of the 22 site objectives. This search leads to a set of metrics that describe the importance of each objective to the rank of the site and a site score for global ranking. The site score depends on the dataset used in the analysis. For this reason, the method is repeated for the brownfield sites, CPPs, and a joint dataset of 250 best brownfields and all CPPs. Given our relatively large dataset of more than +30,000 sites, in the second part, we train a data-driven methodology using Concatenated Neural Networks (ConcNN) that allows predicting the (1) site score, (2) site objectives, and (3) importance value of each objective to the site score. The network needs only the site coordinates and location to make such a prediction.
@@ 3.1 Non-Dominated, Multi-Objective Combinatorial Search
We used a combinatorial search between the 22 objectives to establish a ground-truth outcome through the application of a non-dominated sorting method. When we perform non-dominated sorting across all objectives at once, the resulting Pareto front comprises 1,294 locations, which yields inconclusive results for a single optimal site. To address this limitation, we applied non-dominated sorting to all possible combinations of objectives. A detailed explanation of these combinations is provided in Table 2, which outlines the combinations of objectives for each specified combination length.
@@
Comb. Length Combination Examples Total Number
1 (1), (2), (3), …, (21), (22) 22
2 (1,2), (1,3), (1,4), …, (21,20), (21,22) 231
3 (1,2,3), (1,2,4), (1,2,5), …, (20,21,19), (20,21,22) 1540
21 (1,2, …, 20,21), (1,2, …, 20,22), …, (2,3, …, 21,22) 22
22 (1,2,3, …, 20,21,22) 1
@@
The preprocessed brownfield site objective dataset contains 15,979 rows that represent the individual sites and 22 columns corresponding to objectives. The contiguous U.S. ACRES brownfield dataset includes sites that are often adjacent or in close proximity to one another. Data in the STAND tool rely on state-level and county-level information, proximity to other sites of interest, and integration with geographical databases. We expect locations within close proximity to exhibit identical characteristics. To manage this redundancy, we truncated site coordinates to 0.01 degrees of longitude and latitude, and we removed sites with identical truncated coordinates from the dataset. Later, we inferred the siting results for these removed locations by using results from persisting locations with the same truncated coordinates. This method introduces an uncertainty of approximately 0.315 miles in the southern U.S. and 0.246 miles in the northern U.S. due to the truncation of coordinates. The computational complexity of the non-dominated sorting method varies between $O(N\log N)$ and $O(N^{2})$, depending on the implementation and problem structure. Reducing the number of sites from 34,212 to 15,979 results in a 56.34% decrease in computation time under the low-complexity $O(N\log N)$ scenario and a 77.85% decrease under the high-complexity $O(N^{2})$ scenario.
@@
The non-dominated, multi-objective combinatorial search starts from combination length $s=1$. We only apply non-dominated sorting a single selected objective at a time. For the combination length $s=1$, we apply non-dominated sorting $k\in[1,22]$ times for the examples given in Table 2.
The preprocessed 15,979 rows of data is defined as the matrix $\widehat{A}$ where $\widehat{A}\in\mathbb{R}^{15979\times 22}$. The subsets of $\widehat{A}$ that only include selected column indices are defined as $\widehat{A}_{L_{k}}$, i.e. $\widehat{A}_{L_{k}}\in\widehat{A}$.
@@
We generated the possible combinations of the siting objectives using $C(s,t)$. For $C(1,22)$, there are 22 different $C_{k}$ elements to select, each of these elements is given in Table 2. This operation creates $L$, the possible combinations that choose $s$ elements from a set of $t$ elements regardless of order. $L_{k}$ defines a single combination in the list $L$.
$\begin{split}C(s,22)=L\\ L=\{L_{1},L_{2},L_{3},\ldots,L_{k}\}\\ \end{split}$ (1)
Note that $L$ depends on $s$. For simplicity, we omit the subscript $s$ in the notation for $L$ and the vectors up to $\overrightarrow{SR}$ as discussed in this section.
@@
The function $NS$ is defined as the operation that applies non-dominated sorting on a matrix. The $NS$ function outputs a column vector $\overrightarrow{r}_{k}$ with 15,979 rows indicating the domination for each site. $\overrightarrow{r}_{k}$ has 1 in the ”dominating” $\widehat{A}$ site indices and 0 in the “dominated” site indices. For each $\widehat{A_{L_{k}}}$ subset, there is an unique $\overrightarrow{r}_{k}$ vector.
$NS(\widehat{A}_{L_{k}})=\overrightarrow{r}_{k},\quad\overrightarrow{r}_{k}\in\mathbb{R}^{15979\times 1}$ (2)
The next step involves the summation and normalization of each vector $\overrightarrow{r}_{k}$, resulting in the normalized observation ratio $\overrightarrow{NR}$. Summing the normalized observation ratios across various combination lengths $s$ yields the summed normalized observation ratio $\overrightarrow{SR}$. Finally, applying min-max scaling to $\overrightarrow{SR}$ within the range $[0,1]$ produces $\overrightarrow{\mathbb{M}}$, which acts as a site performance metric for selecting the most suitable reactor sites.
@@
$\begin{split}\overrightarrow{\text{NR}}=\frac{\sum_{k}\overrightarrow{r}_{k}}{\left|\sum_{k}\overrightarrow{r}_{k}\right|}\\ \widehat{SR}=\sum_{s=1}^{22}\widehat{NR}\\ \overrightarrow{\mathbb{M}}=\frac{\overrightarrow{SR}-\min(\overrightarrow{SR})}{\max(\overrightarrow{SR})-\min(\overrightarrow{SR})}\end{split}$ (3)
@@
The methodology presented is based on the non-dominated sorting approach, which we employed to identify the optimal individuals within a given set. Consequently, the siting metric derived from this method is inherently influenced by the specific sites included in the dataset. We demonstrate an example of aquiring the site observation ratio below. In this problem, we select the 3-length combination of the objectives.
$C(3,22)$ $=L=[L_{1},L_{2},L_{3},\ldots,L_{1540}]$
$L_{230}$ $=\begin{bmatrix}2,3,5\end{bmatrix}$
@@
The function $C(s,t)$ creates the list of the combinations. We select the $230^{th}$ item on this listfor demonstration. This lis, $L_{230}$, includes the column (objective) numbers 2, 3, and 5, which represent state electricity price, state electricity imports, and population SVI, respectively. The subset $L_{230}$ of the dataset $\widehat{A}$ shown below:
$\widehat{A}_{L_{230}}$ $=\begin{bmatrix}0.2023&0.6605&0.6158\\ 0.0729&0.2907&0.5883\\ 0.3935&0.5107&0.9067\\ 0.4777&0.7118&0.6409\\ \vdots&\vdots&\vdots\end{bmatrix}$
Applying non-dominated sorting to the dataset subset creates $\overrightarrow{r}_{230}$ as follows:
$NS(\widehat{A}_{L_{230}})$ $=\overrightarrow{r}_{230}=\begin{bmatrix}0\\ 0\\ 1\\ 1\\ \vdots\end{bmatrix}$
@@
This vector indicates that the sites at the first and second row indices are dominated by the other sites in this example. After finding the domination status of all locations for all 3-length combinations of the siting objectives (from 1 to 1,540 total combinations), we normalize the summed vector:
$\overrightarrow{NR}$ $=\begin{bmatrix}0\\ 0\\ 0.0002\\ 0.0002\\ \vdots\end{bmatrix}$
@@
Summing $\overrightarrow{NR}$ for all combination lengths creates the summed normalized observation ratio $\overrightarrow{SR}$, which can then be converted by min-max normalization to $\overrightarrow{\mathbb{M}}$ (the siting metric) as follows:
$\overrightarrow{SR}$ $=\begin{bmatrix}0.013\\ 0.008\\ 0.029\\ 0.034\\ \vdots\end{bmatrix}$
@@
Following the site score, $\overrightarrow{\mathbb{M}}$, we seek to determine which of the 22 objectives contribute the most to the site score. This provides a more informative explanation for a given site’s suitability. To do this, we start by defining the summed normalized objective contributions ($\widehat{SC}$), which show the contribution of each objective to a specific site performance. Then, the dominating indices vector $\overrightarrow{r}_{k}$ is multiplied with $\overrightarrow{v}_{k}^{T}$. The column vector $\overrightarrow{v}_{k}$ shows which objectives are included in the current combination:
@@
$\begin{split}\overrightarrow{v}_{k}[i]=\begin{cases}1&\text{if }i\in L_{k}\\ 0&\text{if }i\notin L_{k}\end{cases}\\ \widehat{R}_{k}=\overrightarrow{r}_{k}\cdot\overrightarrow{v}_{k}^{T}\end{split}$ (4)
Note that $\overrightarrow{v}_{k}$ depends on $s$. For simplicity, the subscript $s$ is omitted in the notation for $\overrightarrow{v}_{k}$ and the matrices up to $\widehat{S}$ as discussed in this section.
@@
The vector $\overrightarrow{v}_{k}$ has 1 at the indices included in the objective combination $L_{k}$ for the combination $k$. The matrix $\widehat{R}_{k}$ resulting from the dot product shows which objectives (column indices) affected the performance of the sites (row indices). We define $\widehat{R}_{k}$ as the contribution matrix for the $k^{th}$ objective combination in the dataset. These contribution matrices are summed for different combinations within the same combination length as follows:
$\widehat{OC}=\sum_{k}\widehat{R}_{k}$

### v2

@@ 2.2 Socioeconomic Characteristics
We extracted nuclear restrictions in each state. Considering it is an uncontrollable parameter for a stakeholder, we define a moratorium as a deal breaker. We express all other restrictions in a single variable as the number of nuclear restrictions in the state. These five remaining restrictions are state legislature approval, state commissioner of environmental protection approval, proving that the construction of a nuclear facility will be economically feasible for ratepayers, demonstrable technology or a means for high-level waste disposal or reprocessing, and proving that the radioactive waste material disposal method will be safe.
We define state electricity price as the annual average electricity price in cents/kWh and electricity imports as the annual amount of electrical power imported by the state in millions of kWh. Our “population sentiment towards nuclear energy” parameter is an average of the 10-year poll findings of sentiment values of the counties in any 20-mile radius. We set state nuclear inclusive policy as a binary value that shows if the state administration has any decisions to support the development of nuclear energy, such as Renewable Portfolio Standards (RPSs), Renewable Portfolio Goals (RPGs), and Clean Energy Standards (CESs).
@@ 2.3 Safety Characteristics
We selected safety objectives as characteristics that support siting a 1000MWt or higher power reactor [22]. Most of the safety objectives (1, 3, 4, 6, 7) are intersections with dangerous areas explained in the list of objectives provided earlier in this section. NRC Regulatory Guide 4.7, “General Site Suitability Criteria for Nuclear Power Stations” describes that siting nuclear reactors close to the publicly used lands depends on local jurisdictions and can be refused. For this reason, the protected land objective checks if the site intersects with American Indian reservations, correctional facilities, critical habitats, forests, hospitals, national monuments, national or state parks, schools or colleges, wild and scenic rivers, wilderness areas, or wildlife refuges. If the site intersects with any of these, the number of intersections is given.
@@ 2.4 Proximity Characteristics
Proximity objectives have been used as binary values as suggested by the STAND tool and previous research conducted in the area. For example, the 10 CFR 100 Regulatory Guide of the U.S. NRC shows that at least a four-mile distance is required between the NPP and the population centers with 25,000 residents [37, 38]. In the STAND tool, this objective is set to 0 if the closest population center distance is closer than four miles. For centers with more than four miles, the objective is linearly scaled between [0,1], where the center with the highest distance from the plant is set to 1.
@@
The retiring power station objective shows if any coal, natural gas, or nuclear power station is in a one-mile radius in the next 20 years. For the sites with a retiring power station nearby, the distances are linearly scaled to the [0,1] range where the lowest-distance site is set to 1. The substation distance objective is the site distance to the closest electricity grid substation in miles, scaled to [0,1] range where the closest site to the substation is set to 1. We define transportation systems as major roads in the United States. The objective is set to 1 if there is a major road closer than 1 mile. For the sites that do not have a major road in a 1-mile radius, the values are linearly scaled to the [0,1] range where the lowest-distance site is set to 1. The streamflow objective shows if there are any freshwater supplies in 20-mile radius, with the minimum 50,000 gallons per minute (GPM) streamflow as recommended in the STAND tool.
The distance values in this section are the distance values that the STAND tool uses for thresholding and linearly scaling the attributes of the selected sites. We selected the generator retirement time (20 years), transportation method (major roads), and streamflow GPM value (50,000) with the guidance of the STAND tool developers.
Lastly, an example of the created dataset is given in Figure 1. This dataset is scaled to the [0,1] range afterwards, higher values for each attribute indicate favorable objectives. For example, 0.9 in the state electricity imports column indicates a high electricity import in the state, while 0.1 indicates low electricity import.
@@ 2.5 Coal Power Plant Sites
According to the U.S. Department of Energy, there are currently 157 inactive and 237 active candidate coal power plant locations that are suitable for housing NPPs [13]. These sites are accompanied by pertinent data, including the 22 objectives outlined in the STAND framework, which are detailed in previous sections and in Table 1.
Previous research identified 265 U.S. coal power plant sites as relevant for NPP siting [22]. Additionally, the coal-to-nuclear transition dataset was updated in October 2024 by the Fastest Path to Zero Initiative at the University of Michigan. We have adapted the coal power plant dataset, which was previously examined using the weighted sum method, for multi-objective optimization. By using this dataset, we facilitated a one-to-one comparative analysis with the Brownfield sites.
The coal-to-nuclear siting dataset is considerably smaller than the Brownfield siting dataset, therefore the addition of the CPP dataset to the Brownfield dataset does not significantly increase the processing time. We have created a joint dataset with Brownfields and CPPs to analyze and compare them. The joint dataset has then been used to train the machine learning models described later in the next section. Processing all sites as a single dataset allowed a comparison between the optimal locations identified in both datasets.
@@ 3 Research Methodology
This section describes the methodology used for the assessment of the Brownfield and coal sites described in Section 2, which consists of two major parts. First, we describe our high-dimensional non-dominated combinatorial search approach that is used to rank all sites based on how many times they show up in the Pareto-front when considering every possible combination of the 22 site objectives. This search leads to a set of metrics that describe the importance of each objective to the rank of the site and a site score for global ranking. The methodology is applied to the joint dataset of 34211 Brownfield sites and 265 CPP sites. Given our relatively large dataset of more than +30,000 sites, in the second part, we train a data-driven methodology using Concatenated Neural Networks (ConcNN) and lookup tables that allow predicting the (1) site score, (2) site objectives, and (3) importance value of each objective to the site score. The network needs only the site location to make such a prediction.
The proposed methodology in this section provides a comprehensive, unbiased framework for aggregating results across all possible combinations of objectives, avoiding analyst-defined weights that might introduce subjective bias. This exhaustive approach ensures that the siting metric reflects a diverse range of objective trade-offs, which is particularly useful in high-dimensional decision-making problems. By leveraging non-dominated sorting for each subset of objectives, the method captures multi-objective interactions that might otherwise be overlooked, enhancing the robustness of the analysis.
The siting metric described in section 3.1 provides a straightforward assessment of each site’s performance relative to others in the dataset. Essentially, it functions as a dominance metric, quantifying the extent to which a site’s objectives outperform those of other sites. This metric is purely mathematical, relying on the non-dominated sorting method, a widely used approach in multi-objective optimization problems.
@@ 3.1 Non-Dominated, Multi-Objective Combinatorial Search
We used a combinatorial search between the 22 objectives to ensure a bias-free outcome through the application of a non-dominated sorting method. When we perform non-dominated sorting across all objectives at once, the resulting Pareto front comprises 1,294 locations, which yields inconclusive results for a single optimal site. To address this limitation, we applied non-dominated sorting to all possible combinations of objectives. A detailed explanation of these combinations is provided in Table 2, which outlines the combinations of objectives for each specified combination length.
@@
Combination Length ($s$) Combination Examples ($L_{s,k}$) Total Number (N)
1 (1), (2), (3), $\dots$, (21), (22) 22
2 (1,2), (1,3), (1,4), $\dots$, (21,20), (21,22) 231
3 (1,2,3), (1,2,4), $\dots$, (20,21,19), (20,21,22) 1,540
4 (1,2,3,4), (1,2,3,5), $\dots$, (19,20,21,22) 7315
5 (1,2,3,4,5), (1,2,3,4,6), $\dots$, (18,19,20,21,22) 26,334
6 (1,2, $\dots$, 5,6), $\dots$, (17,18, $\dots$, 21,22) 74,613
7 (1,2, $\dots$, 6,7), $\dots$, (16,17, $\dots$, 21,22) 170,544
8 (1,2, $\dots$, 7,8), $\dots$, (15,16, $\dots$, 21,22) 319,770
9 (1,2, $\dots$, 8,9), $\dots$, (14,15, $\dots$, 21,22) 497,420
10 (1,2, $\dots$, 9,10), $\dots$, (13,14, $\dots$, 21,22) 646,646
11 (1,2, $\dots$, 10,11), $\dots$, (12,13, $\dots$, 21,22) 705,905
12 (1,2, $\dots$, 11,12), $\dots$, (11,12, $\dots$, 21,22) 646,646
13 (1,2, $\dots$, 12,13), $\dots$, (10,11, $\dots$, 21,22) 497,420
14 (1,2, $\dots$, 13,14), $\dots$, (9,10, $\dots$, 21,22) 319,770
15 (1,2, $\dots$, 14,15), $\dots$, (8,9, $\dots$, 21,22) 170,544
16 (1,2, $\dots$, 15,16), $\dots$, (7,8, $\dots$, 21,22) 74,613
17 (1,2, $\dots$, 16,17), $\dots$, (6,7, $\dots$, 21,22) 26,334
18 (1,2, $\dots$, 17,18), $\dots$, (5,6, $\dots$, 21,22) 7,315
19 (1,2, $\dots$, 18,19), $\dots$, (4,5, $\dots$, 21,22) 1,540
20 (1,2, $\dots$, 19,20), $\dots$, (3,4, $\dots$, 21,22) 231
21 (1,2, $\dots$, 20,21), $\dots$, (2,3, $\dots$, 21,22) 22
22 (1, 2, 3, $\dots$, 20, 21, 22) 1
Total Combination Sum: 4,194,303
@@
The preprocessed joint dataset (both Brownfield and coal sites) contains 16,057 rows that represent the individual sites and 22 columns corresponding to objectives. The contiguous U.S. ACRES Brownfield dataset includes sites that are often adjacent or in close proximity to one another. The data in the STAND tool rely on state-level and county-level information, proximity to other sites of interest, and integration with geographical databases. We expect locations within close proximity to exhibit identical characteristics. To manage this redundancy, we truncated site coordinates to 0.01 degrees of longitude and latitude, and we removed sites with identical truncated coordinates from the dataset. Later, we inferred the siting results for these removed locations by using results from persisting locations with the same truncated coordinates. This method introduces an uncertainty of approximately 0.315 miles in the southern U.S. and 0.246 miles in the northern U.S. due to the truncation of coordinates. The computational complexity of the non-dominated sorting method varies between $O(N\log N)$ and $O(N^{2})$, depending on the implementation and problem structure. Reducing the number of total sites from 34,476 to 16,057 results in a 64.23% decrease in computation time under the low-complexity $O(N\log N)$ scenario and a 77.85% decrease under the high-complexity $O(N^{2})$ scenario.
@@
The non-dominated, multi-objective combinatorial search starts from combination length $s=1$. We only apply non-dominated sorting to a single selected objective at a time. For the combination length $s=1$, we apply non-dominated sorting $k\in[1,N=22]$ times for the examples given in Table 2. For the combination length $s=2$, we apply non-dominated sorting $k\in[1,N=231]$ times, and so on.
The preprocessed dataset consists of 16,057 rows, represented as the matrix $\widehat{A}\in\mathbb{R}^{16057\times 22}$. Subsets of $\widehat{A}$ that include only selected column indices are denoted as $\widehat{A}_{L_{s,k}}$, also referred to as $\widehat{B}_{s,k}$. Each $\widehat{B}_{s,k}$ matrix corresponds to a specific combination of selected columns, indexed by $s$ and $k$. To encapsulate all possible subsets, we construct a single tensor $\widehat{B}$ that consolidates this subset information.
@@
We generated the possible combinations of the siting objectives using the combinatorial function $C(t,s)$. For $C(22,1)$, there are 22 different $L_{s,k}$ elements to select, each of these elements is given in Table 2. This operation creates $L$, the list of possible combinations that choose $s$ elements from a set of $t$ elements regardless of order. $L$ is a list of ordered index sets. $L_{s,k}$ defines a single combination in the list $L$, where $L_{1,1}=(1)$, $L_{1,2}=(2)$, $L_{2,1}=(1,2)$ which is shown as:
$\begin{split}C(22,s)=L\\ L=\{L_{s,1},L_{s,2},L_{s,3},\ldots,L_{s,N}\}\\ \text{where N is the total number of combinations for s.}\\ \end{split}$ (1)
@@
The function $NS$ is defined as the operation that applies non-dominated sorting on a matrix or tensor. In Eq.(2), $i$ represents the site index and $j$ represents the objectives of the selected subset ($\widehat{B}_{s,k}$). The NS function is applied along the site index ($i$) axis for each subset defined by the ($s,k$) pair. The $NS$ function outputs a tensor $\widehat{r}$ with 16,057 length in $i$ axis indicating the domination for each site in the selected subset. For a given subset $\widehat{B}_{s,k}$, $r_{s,k,i}$ is 1 for the “dominating” $\widehat{B}$ site indices ($i$) and 0 for the “dominated” site indices. For each matrix slice of $\widehat{B}$ corresponding to a given $(s,k)$ pair, there exists a unique vector slice of the $\widehat{r}$ tensor that represents the results associated with that specific $(s,k)$ pair.
$\widehat{B}_{s,k}=\widehat{A}_{L_{s,k}}\quad,\quad\widehat{B}\in\mathbb{R}^{22\times N\times 16057\times 22}\quad,\quad\text{with indices }(s,k,i,j)$ (2)
$NS(\widehat{B})=\widehat{r}\quad,\quad\widehat{r}\in\mathbb{R}^{22\times N\times 16057}\quad,\quad\text{with indices }(s,k,i).$
The next step involves the normalization of the tensor $\widehat{r}$ on its $k$ axis. Normalized $\widehat{r}$ tensor is shown as $\widehat{r^{\prime}}$. The sum of $\widehat{r^{\prime}}$ along the $k$ axis produces the normalized observation ratio matrix $\widehat{NR}$.
@@
$\begin{split}r^{\prime}_{s,k,i}=\frac{r_{s,k,i}}{\sum_{k}r_{s,k,i}}\quad,\quad\forall s,i\\ \widehat{r^{\prime}}\in\mathbb{R}^{22\times N\times 16057}\quad,\quad\text{with indices }(s,k,i)\\ NR_{s,i}=\sum_{k}r^{\prime}_{s,k,i}\quad,\quad\widehat{NR}\in\mathbb{R}^{22\times 16057}\quad,\quad\text{with indices }(s,i).\\ \end{split}$ (3)
The final summed normalized observation ratio $\overrightarrow{SR}$ is obtained by summing $\widehat{NR}$ over all combination lengths $s$. Next, we apply min-max scaling to $\overrightarrow{SR}$ within the range $[0,1]$, which produces $\overrightarrow{\mathbb{M}}$ that acts as a site performance metric for selecting the most suitable reactor sites. These two variables are defined as follows:
$\begin{split}SR_{i}=\sum_{s=1}^{22}NR_{s,i}\quad,\quad\overrightarrow{SR}\in\mathbb{R}^{16057}\quad,\quad\text{with index }(i)\\ \overrightarrow{\mathbb{M}}=\frac{\overrightarrow{SR}-\min(\overrightarrow{SR})}{\max(\overrightarrow{SR})-\min(\overrightarrow{SR})}\quad,\quad\overrightarrow{\mathbb{M}}\in\mathbb{R}^{16057}\quad,\quad\text{with index }(i).\end{split}$ (4)
@@
The methodology presented is based on the non-dominated sorting approach, which we employed to identify the optimal individuals within a given set. Consequently, the siting metric derived from this method is inherently influenced by the specific sites included in the dataset. We demonstrate an example of acquiring the site observation ratio below. In this example, we select the 3-length combination of the objectives.
$C(22,3)=L[s=3,k=230]=[L_{3,1},L_{3,2},L_{3,3},\ldots,L_{3,1540}]$
[…]
@@
The function $C(s,t)$ creates the list of the combinations. We arbitrarily select the $230^{th}$ item on this list for demonstration. This list, $L_{s,k}[s=3,k=230]$, includes the column (objective) numbers 2, 3, and 5, which represent state electricity price, state electricity imports, and population SVI, respectively. The subset $\widehat{A}_{L_{s,k}}[s=3,k=230]$ of the dataset $\widehat{A}$ (or tensor $\hat{B}$ as defined before) is shown below:
$\widehat{B}[s=3,k=230]$ $=\begin{bmatrix}0.2023&0.6605&0.6158\\ 0.0729&0.2907&0.5883\\ 0.3935&0.5107&0.9067\\ 0.4777&0.7118&0.6409\\ \vdots&\vdots&\vdots\end{bmatrix}$
The subset $\widehat{B}[s=3,k=230]$ of $\widehat{B}$ shows a single matrix slice in the 4D $\widehat{B}$ tensor. Applying non-dominated sorting to the dataset creates $\widehat{r}$ tensor. Fixing the $s$ and $k$ variables and leaving the site index ($i$) variable free in this 3D tensor results in its vector slice $\widehat{r}[s=3,k=230]$:
$NS(\widehat{B}[s=3,k=230])$ $=\widehat{r}[s=3,k=230]=\begin{bmatrix}0\\ 0\\ 1\\ 1\\ \vdots\end{bmatrix}$
@@
This vector indicates that the sites at the first and second row indices are dominated by the other sites in this example. After finding the domination status of all locations for all 3-length combinations of the siting objectives (from 1 to 1,540 total combinations), we normalize the summed vector and sum for all k values. These operations create $\widehat{NR}$. Fixing the $s=3$ and $k$ variables and leaving the site index ($i$) variable free creates the vector $\widehat{NR}[s=3]$:
$\widehat{NR}[s=3]$ $=\begin{bmatrix}0\\ 0\\ 0.0002\\ 0.0002\\ \vdots\end{bmatrix}$
@@
Summing $\widehat{NR}$ for all combination lengths ($s$) creates the summed normalized observation ratio $\overrightarrow{SR}$, which can then be converted by min-max normalization to $\overrightarrow{\mathbb{M}}$ (the siting metric) as follows:
$\widehat{SR}$ $=\begin{bmatrix}0.013\\ 0.008\\ 0.029\\ 0.034\\ \vdots\end{bmatrix}$
@@
Following the site score, $\overrightarrow{\mathbb{M}}$, we seek to determine which of the 22 objectives contribute the most to the site score. This provides a more informative explanation for a given site’s suitability. This information is shown in the summed normalized objective contributions matrix ($\widehat{SC}$), which shows the contribution of each objective (j) to a specific site performance (i). To gather this, we start by defining a slice of the dominating indices tensor $\widehat{r}$ from the previous derivation. This slice is named $\widehat{r}_{s,k}$ as a column vector. $\widehat{r}_{s,k}$ is multiplied with $\overrightarrow{v}_{s,k}^{T}$ row vector. The row vector $\overrightarrow{v}_{s,k}^{T}$ shows which objectives are included in the current combination:
@@
$\begin{split}\overrightarrow{v}_{s,k}[j]=\begin{cases}1&\text{if }j\in L_{s,k}\\ 0&\text{if }j\notin L_{s,k}\end{cases}\quad,\quad v_{s,k}\in\mathbb{R}^{22\times 1}\quad,\quad r_{s,k}\in\mathbb{R}^{16057\times 1}\\ \widehat{R}_{s,k}=\widehat{r}_{s,k}\cdot\overrightarrow{v}_{s,k}^{T}\quad,\quad\widehat{R}\in\mathbb{R}^{22\times N\times 16057\times 22}\quad,\quad\text{with indices }(s,k,i,j).\end{split}$ (5)
@@
The vector $\overrightarrow{v}_{s,k}^{T}$ has 1 at the indices included in the objective combination $L_{s,k}$ for the combination $k$. The 4D tensor $\widehat{R}$ results from merging the matrices for different $s$ and $k$ sets for each dot product result. $\widehat{R}$ shows which objectives ($j$) affected the performance of the sites ($i$). We define $\widehat{R}$ as the contribution matrix for $s^{th}$ combination length and the $k^{th}$ objective combination in the dataset. These contribution matrices are summed for different combinations within the same combination length as follows:
$OC_{s,i,j}=\sum_{k}R_{s,k,i,j}\quad,\quad\widehat{OC}\in\mathbb{R}^{22\times 16057\times 22}.$ (6)